For a reaction of conversion of ATP into ADP at 20⁰C, calculate the reaction Gibbs free energy. Given the change in enthalpy is 20.27 Kcal and the change in entropy is 85 cal/Kelvin


1
Expert's answer
2022-01-24T18:44:37-0500

To be given in question

T=273+30=293K

∆H=20.27Kcal

∆S=85cal/K

We know that

"\u2206G=\u2206H-T\u2206S"

Put value


"∆G=20270-293\times85=20270-24905=-4635Cal"

"\u2206G=-4.635Kcal"